# Prompting Large Language Models with Divide-and-Conquer Program for Discerning Problem Solving

### 14 Mar 2024 v3

@@ Abstract
Foundation models, such as Large language Models (LLMs), have attracted significant amount of interest due to their large number of applications. Existing works show that appropriate prompt design, such as Chain-of-Thoughts, can unlock LLM’s powerful capacity in diverse areas. However, when handling tasks involving repetitive sub-tasks and/or deceptive contents, such as arithmetic calculation and article-level fake news detection, existing prompting strategies either suffers from insufficient expressive power or intermediate errors triggered by hallucination. To make LLM more discerning to such intermediate errors, we propose to guide LLM with a Divide-and-Conquer program that simultaneously ensures superior expressive power and disentangles task decomposition, sub-task resolution, and resolution assembly process. Theoretic analysis reveals that our strategy can guide LLM to extend the expressive power of fixed-depth Transformer. Experiments indicate that our proposed method can achieve better performance than typical prompting strategies in tasks bothered by intermediate errors and deceptive contents, such as large integer multiplication, hallucination detection and misinformation detection.
@@ 1 Introduction
In fact, human brains also suffer from similar hallucination issues (Liu & Sajda, 2023), especially when the tasks are too hard or too complex. For example, when reviewing a long academic paper, some reviewers produce low-quality reviews (Garcia et al., 2021; Tennant & Ross-Hellauer, 2020; Cortes & Lawrence, 2021) containing hallucination-like intermediate errors, such as pointing out some ‘missing baselines’ that have already been sufficiently discussed by authors and requiring the authors to conduct ablation studies for non-existent sub-modules. To avoid such ridiculous mistakes, experienced reviewers usually think slowly (Kahneman, 2011) to follow a Divide-and-Conquer paradigm to handle this task. Specifically, they decompose the paper review as examinations of multiple central opinions. Then they retrieve supporting corpus to verify them respectively. Similar strategies are also applied in automated evaluation to the quality of text generated by LLMs (Cui et al., 2023).
@@
To validate the expressive power of our proposed prompting strategy, we provide a theoretic analysis to validate that the proposed strategy can expand the expressive power of fixed-depth log-precision Transformer. To further empirically validate the advantage of our proposed method, we evaluate our method and representative baselines on three tasks that are challenging to existing prompting strategies even on state-of-the-art LLMs: Large Integer Multiplication, Hallucination Detection, Article-level Fact Verification (Cheng & Zhang, 2023; Li et al., 2023a; Wadden et al., 2020; Hu et al., 2023; Wu et al., 2023). These tasks either require very long reasoning paths (e.g. large integer multiplication) or contain deceptive contents (e.g. hallucination detection and fact verification), making existing methods like Chain-of-Thought prompting prone to missing intermediate steps. Meanwhile, these tasks share a common nature that the correctness of the final answer substantially relies on the thoroughness of all steps. Our experimental results show that the proposed method outperforms the baselines on all three tasks.
@@ 2.2 Prompting Strategies for Large Language Models
The aforementioned CoT and EoT families incorporate LLM with stronger expressive power than IO prompting. However, a critical issue of them is that, they could miss or ignore some important intermediate steps or contents (Liu et al., 2023). This problem is even worse when we are handling tasks involved with long input (e.g. long documents and large numbers). Typical examples include large number arithmetic calculation and fact verification in long documents. Compared to them, Least-to-Most prompting and Decomposed Prompting introduce explicit task decomposition to enumerate sub-tasks. However, their task decomposers are based on multi-round conversation or question-answering, which often guides the LLM to sequentially tackle the sub-tasks. When tackling contents containing deceptive contents, e.g. hallucination and fake news detection, such design has two issues. First, it will navigate the LLM through the deceptive content’s flow and thus make LLM prone to deception. Second, as pointed by previous works (Yang & Ettinger, 2023; Dziri et al., 2024), when the round of conversation increases, the hallucination/deception will propagate and accumulate in the contexts, leading to decreased performance and accuracy.
@@ 4.1 Fixed-Depth Transformer can not solve 2-color Binary Subtree Isomorphism (2-BSI)
4.2 Fixed-Depth Transformer guided by Divide-and-Conquer program solves 2-BSI
@@
There exists a log-precision transformer with fixed depth $L$ and hidden dimension $d$ that can solve the 2-BSI of any size with fixed-length prompt $m$ (for merge), $t$ (for sub-task tackling) and $d$ (for task decomposition).
Proof Sketch: The detailed proof to this theorem is provided in the Appendix A.1. Here we give a brief flow of the proof. To prove this theorem, we first show that there exists an problem size function $f(\cdot)$, hyper-parameter $w$, merge function $m(\cdot)$, sub-task tackling function $t(\cdot)$ and decomposition function $d(\cdot)$ that can solve the problem with divide-and-conquer strategy. Then we will prove that there exists one log-precision transformer with fixed depth $L$ and hidden dimension $d$ that can express $a(\cdot)$, $t(\cdot)$ and $d(\cdot)$ respectively with different but fixed-length prompts. In this way, we can prove the theorem.

### 05 Apr 2024 v4

# Prompting with Divide-and-Conquer Program Makes Large Language Models Discerning to Hallucination and Deception

@@ Abstract
Foundation models, such as Large language Models (LLMs), have attracted significant amount of interest due to their large number of applications. Existing works show that appropriate prompt design, such as Chain-of-Thoughts, can unlock LLM’s powerful capacity in diverse areas. However, when handling tasks involving repetitive sub-tasks and/or deceptive contents, such as arithmetic calculation and article-level fake news detection, existing prompting strategies either suffers from insufficient expressive power or intermediate errors triggered by hallucination. To make LLM more discerning to such intermediate errors, we propose to guide LLM with a Divide-and-Conquer program that simultaneously ensures superior expressive power and disentangles task decomposition, sub-task resolution, and resolution assembly process. Theoretic analysis reveals that our strategy can guide LLM to extend the expressive power of fixed-depth Transformer while exploiting smaller context window size compared to existing prompting strategies like CoT. Experiments indicate that our proposed method can achieve better performance than typical prompting strategies in tasks bothered by intermediate errors and deceptive contents, such as large integer multiplication, hallucination detection and misinformation detection.
@@ 1 Introduction
In fact, human brains also suffer from similar hallucination issues (Liu & Sajda, 2023), especially when the tasks are too hard or too complex. For example, when reviewing a long academic paper, some reviewers produce low-quality reviews (Garcia et al., 2021; Tennant & Ross-Hellauer, 2020; Cortes & Lawrence, 2021) containing hallucination-like intermediate errors, such as pointing out some ‘missing baselines’ that have already been sufficiently discussed by authors and requiring the authors to conduct ablation studies for non-existent sub-modules. To avoid such ridiculous mistakes, experienced reviewers usually think slowly (Kahneman, 2011) to follow a Divide-and-Conquer paradigm to handle this task. Specifically, they decompose the paper review as examinations of multiple central opinions. Then they retrieve supporting corpus to verify them respectively. Similar strategies are also applied in decoding to specifically improve LLM’s capacity in math (Lee & Kim, 2023) and automated evaluation (Cui et al., 2023).
@@
To validate the expressive power of our proposed prompting strategy, we provide a theoretic analysis to validate that the proposed strategy can expand the expressive power of fixed-depth log-precision Transformer. Meanwhile, the theoretic analysis to computational complexity also reveals that compared to sequential prompting methods like CoT and LtM which can also extend the expressive power, our proposed method generally require smaller context window size for the Transformer (i.e., $O(|t_{b}|)$ vs $O(|t_{p}|\times|t_{b}|)$), alleviating the error propagation in long context (Yang & Ettinger, 2023). To further empirically validate the advantage of our proposed method, we evaluate our method and representative baselines on three tasks that are challenging to existing prompting strategies even on state-of-the-art LLMs: Large Integer Multiplication, Hallucination Detection, Article-level Fact Verification (Cheng & Zhang, 2023; Li et al., 2023a; Wadden et al., 2020; Hu et al., 2023; Wu et al., 2023). These tasks either require very long reasoning paths (e.g. large integer multiplication) or contain deceptive contents (e.g. hallucination detection and fact verification), making existing methods like Chain-of-Thought prompting prone to missing intermediate steps. Meanwhile, these tasks share a common nature that the correctness of the final answer substantially relies on the thoroughness of all steps. Our experimental results show that the proposed method outperforms the baselines on all three tasks.
@@ 2.2 Prompting Strategies for Large Language Models
The aforementioned CoT and EoT families incorporate LLM with stronger expressive power than IO prompting. However, a critical issue of them is that, they could miss or ignore some important intermediate steps or contents (Liu et al., 2023). This problem is even worse when we are handling tasks involved with long input (e.g. long documents and large numbers). Typical examples include large number arithmetic calculation and fact verification in long documents. Compared to them, Least-to-Most prompting and Decomposed Prompting introduce explicit task decomposition to enumerate sub-tasks. However, their task decomposers are based on multi-round conversation or question-answering, which often guides the LLM to sequentially tackle the sub-tasks. When tackling contents containing deceptive contents, e.g. hallucination and fake news detection, such design has two issues. First, it will navigate the LLM through the deceptive content’s flow and thus make LLM prone to deception. Second, as pointed by previous works (Yang & Ettinger, 2023; Dziri et al., 2024), when the round of conversation increases, the context window size for auto-regressive decoding also increases, making the hallucination/deception propagate and accumulate in the contexts, leading to decreased performance and accuracy.
@@ 4.1 Fixed-Depth Transformer can not solve 2-color Binary Subtree Isomorphism (2-BSI)
4.2 Divide-and-Conquer program guide Fixed-Depth Transformer to solve 2-BSI better than IO-prompting and Sequential Prompting
@@
With Divide-and-Conquer prompting program, a log-precision transformer with fixed depth $L$ and fixed hidden dimension $d$ can solve the 2-BSI of any size with fixed-length prompt $m$ (for merge), $t$ (for sub-task tackling) and $d$ (for task decomposition). The context length and deconding length for each calling is $O(|t_{b}|)$ and the DaC prompting program will call the transformer $O(|t_{p}|)$ times.
Proof Sketch: The detailed proof to this theorem is provided in the Appendix A.1. Here we give a brief flow of the proof. To prove this theorem, we first show that there exists an problem size function $f(\cdot)$, hyper-parameter $w$, merge function $m(\cdot)$, sub-task tackling function $t(\cdot)$ and decomposition function $d(\cdot)$ that can solve the problem under the guidance of a divide-and-conquer program. The input length and output length of every aforementioned function are both $O(|t_{b}|)$, and the divide-and-conquer program will call them $O(|t_{p}|)$ times in total. Then we will prove that there exists one log-precision transformer with fixed depth $L$ and hidden dimension $d$ that can express $a(\cdot)$, $t(\cdot)$ and $d(\cdot)$ respectively with different but fixed-length prompts. In this way, we can prove the theorem.
The above theorem explains why DaC program achieves better expressive power compared to IO-prompting. Meanwhile, from the above theoretic results, we can acquire a corollary:
Corollary 4.3.
A sequential prompting strategy (e.g., CoT and LtM) can simulate the trace of DaC prompting strategy to solve the 2-BSI problem. However, their average context length during decoding is $O(|t_{p}|\times|t_{b}|)$.
The above corollary reveals that when applying sequential prompting strategy such CoT and LtM to solve the problems with same algorithms like DaC, although both sequential prompting strategy and DaC prompting share the same decoding length, the maximum context length they require are different. Since DaC program follows the disentangled-sub-process principle, a transformer with context length of $O(t_{b})$ is sufficient. But for sequential prompting strategies, their maximum context length during decoding will be $O(|t_{p}|\times|t_{b}|)$. As claimed by (Yang & Ettinger, 2023), long context length during decoding will lead to the propagation of errors and hallucinations in the reasoning path.